OpenStudy (anonymous):

The length of a rectangle is 5 meters more than its width. If the area is 84 square meters, what are the length and width?

OpenStudy (anonymous):

x(x+5) =84 ===> x^2+5x-84=0 ===> (x-7)(x+12)=0 ===> x=7 or x=-12

OpenStudy (anonymous):

area of rectangle=length*width length=5+width so area=(5+width)*width assume that width=w; area=(5+w)*w 84=5w+w^2 w^2+5w-84=0 (w+12)*(w-7)=0 w=7 w=-12 however length can not be negative so w=7 and length=12

OpenStudy (anonymous):

so width = 7 ==> lenght = 12

OpenStudy (anonymous):

yes